OpenStudy (anonymous):

Express the required measurement in exact form. A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground? help me please

OpenStudy (anonymous):

circumference of wheel = 2*26*pi length of the arc = (80/360) * 2*26*pi now u can use calculator to find this it comes to 36.3"

OpenStudy (anonymous):

oh - in exact form it is (2/9)52pi = 104pi/9